OpenStudy (anonymous):

True or false: The equation tan2x + 1 = sec2x.

OpenStudy (jziggy):

Is this tan(2x) + 1 = sec(2x) or tan^2(x) + 1 = sec^2(x)?

OpenStudy (anonymous):

Tan^2x + 1 = sec^2x

OpenStudy (jziggy):

Start with the basic Pythagorean identity sin^2(x) + cos^2(x) = 1 and then divide the entire equation by cos^2(x)

OpenStudy (anonymous):

Okay thanks